OpenStudy (anonymous):

A single mass m1 = 3.1 kg hangs from a spring in a motionless elevator. The spring is extended x = 11 cm from its unstretched length. 1)What is the spring constant of the spring?

OpenStudy (stormfire1):

F = mg so F = 3.1kg * 9.8m/s^2 => 30.38 kg*m/s^2 = 30.38N Now calculate the spring constant using the formula F = -kx. Rearranging things you get k = -F/x...so k = -30.38N/.011m k = 2761.82

OpenStudy (stormfire1):

and that's in N*m (newton meters)

OpenStudy (anonymous):

So it is only asking for magnitude?

OpenStudy (anonymous):

Isn't it 276.182?

OpenStudy (stormfire1):

oops...you're right. I inadvertently used .011m when it should have been .11m

OpenStudy (stormfire1):

that answer sounds more like a k-value too...the first answer sounded a bit high